Systems Staff Meeting Notes 03/08/04

Systems Staff Meeting Notes
March 8, 2004

Present: Cynthia Bowker, Will Chaney, Bill Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
  1. It was a beautiful weekend.
  2. Ying took Selina to Sea World. She was more interested in the dogs than the sea creatures.
  3. Geoffrey went out into the woods this weekend. Lots of pollen, flowers blooming. Lots of leaves falling as well.
  4. TLC Web site – if you vote, you have a 1 in 3 chance of having your mortgage paid off.
Moves/Construction Status
  1. According to Lori’s latest email, by the end of today all of the various moves to ALF will be complete and accessible. The number of volume requests should be increasing.
  2. West is now decommissioned. The building gets turned over tomorrow to the construction company. There is no reason for us to be over there anymore. We will not be telling them to unplug at the core. We’ve notified CNS that all of the network equipment was removed. They will figure it out by the lack of traffic. 238 & 239 subnets are split to Journalism and Education still – may be done by magic at the core router.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Orders – 50 headless machines on order. The configuration changed – could not get the skew for inside delivery with the other machines. These are bigger in size, but have the same performance and are much cheaper. No distribution list available yet.
  2. The Dell rep also offered us a 2-for-1 printer special: 35 pages per minute, 175,000 cycle printer. One will be for ALF, one for the shelf.
  3. Barcode Readers – we are currently buying the best possible, but now we have a request From Resource Services for something even better – 5 wands, single point type.
  4. With the Library in its new service mode and since patrons are scarce with the closing of West, we’ve had a request for a weatherproof laptop and laser printer that can be used by a roving librarian out in the plaza and elsewhere around campus. This set-up is to be mounted on a popcorn cart, so as to be sturdy and attract business. Added suggestions from Systems include: solar panels, a calliope and a monkey.
  1. Spam is getting so good now – it is so well-composed – a recent one says you’ve charged your credit card for goods and tells you to click a link to check on it – this type of spam has about a 90% hit rate.
  1. Power Outage – south tower was not affected. 8:00-10:00 AM did not happen. The outage ran longer than two hours. The North tower was shut down. We shut down the server room, pulled conditioned power, all UPS, phone system, three servers in Bill’s office and the north tower. Presumably, everything is all plugged back in with no problems. There was a glitch with the Fenwall system in the server room – it was alerting – seems to have lost the house current. It is not on the conditioned power – may be on the north tower power. They only put in disconnect, which means we will probably have another power outage soon.
  2. Server Room – temperature got to 70 degrees on Friday. The new AC was putting out hot air. The breaker was tripped (it is located in the old server room), so Bill reset it. Sunday morning, the breaker was tripped again and it was blowing hot air. Workmen are currently fixing the problem.
  3. Window project is way overdue.
  4. CNS will install wireless on the 4th floor of MSL – they’re hoping it will bleed through to the 3rd & 5th floors. Supposed to go in over Spring Break.
  5. Old WSID cleanup – mostly done, had one machine in LAC today – it’s now back online.
  6. User ID cleanup – complete, more killed than remained.
  7. AD meeting next week – will be reporting on the status of AD on campus.
  8. New machines – will be staff replacements. They will have DVD burners, but not the software to run it. Need to get a machine back for each one we give out.
Software Projects Status
  1. Ref Express ASP page provisions – extreme programming project to start today. Same dynamic duo (Cynthia and Hilton) will be at the helm.
  2. Firefox ongoing – Cynthia reports that NetG pop-ups were blocked. Firefox will allow you to list sites that are exempt from pop-ups.
  3. Maps/GIS – Joe has been camped out on the server console quite a bit. He updated the database with city info, after that, it did not work. Data is stored in upper case. Need to standardize the interface. Need to contact Arch IMS to see if that function can be changed.
Web Projects
  1. Geoffrey has been made aware of a number of observations that the Systems Web page does not conform to the Library standards. There is not currently a Library standard as of yet, which is why we have complied.

Back to the Systems Homepage

Last updated March 15, 2004
by Debra Fetzer